7 is 31% of what total amount
Found 2 solutions by schrammbledeggs, stanbon:
Answer by schrammbledeggs(41)

Algebraically written...this can be modeled by
0.31x = 7
Divide both sides by 0.31
x = 7 / 0.31
x = 22.58 (approximately)

Answer by stanbon(75887)

7 is 31% of what total amount
------------
EQUATION:
7 = 0.31x
x = 7/0.31 = 22.58...
Cheers,
Stan H.